# Improving Student-AI Interaction Through Pedagogical Prompting: An Example in Computer Science Education

### v1

@@
Abstract.
With the proliferation of large language model (LLM) applications since 2022, their use in education has sparked both excitement and concern. Recent studies consistently highlight students’ (mis)use of LLMs can hinder learning outcomes. This work aims to teach students how to effectively prompt LLMs to improve their learning. We first proposed pedagogical prompting, a theoretically-grounded new concept to elicit learning-oriented responses from LLMs. To move from concept design to a proof-of-concept learning intervention in real educational settings, we selected early undergraduate CS education (CS1/CS2) as the example context. We began with a formative survey study with instructors (N=36) teaching early-stage undergraduate-level CS courses to inform the instructional design based on classroom needs. Based on their insights, we designed and developed a learning intervention through an interactive system with scenario-based instruction to train pedagogical prompting skills. Finally, we evaluated its instructional effectiveness through a user study with CS novice students (N=22) using pre/post-tests. Through mixed methods analyses, our results indicate significant improvements in learners’ LLM-based pedagogical help-seeking skills, along with positive attitudes toward the system and increased willingness to use pedagogical prompts in the future. Our contributions include (1) a theoretical framework of pedagogical prompting; (2) empirical insights into current instructor attitudes toward pedagogical prompting; and (3) a learning intervention design with an interactive learning tool and scenario-based instruction leading to promising results on teaching LLM-based help-seeking. Our approach is scalable for broader implementation in classrooms and has the potential to be integrated into tools like ChatGPT as an on-boarding experience to encourage learning-oriented use of generative AI.
Prompt Engineering, Human-AI Interaction, Large-language Model, Pedagogical Environments, AI-Assisted Learning, Educational Technology, CS1, Programming Novices, Learning Sciences
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Applied computing Interactive learning environmentsccs: Applied computing Computer-assisted instructionccs: Human-centered computing Interactive systems and toolsccs: Social and professional topics Computing education
1. Introduction
@@
To promote more learning-oriented use of LLMs, researchers have proposed tools that embed pedagogical guardrails (Liffiton et al., 2023; Kazemitabaar et al., 2024; Xiao et al., 2024a; Hou et al., 2024). These tools aim to provide programming support without revealing direct solutions (Pirzado et al., 2024). While such designs encourage deeper engagement, they often constrain the full capabilities of LLMs and fall short of aligning with learners’ expectations or real-world needs. As a result, learners tend to underutilize these systems, favoring more versatile tools like ChatGPT, even if these are used in ways that hinder long-term learning (Kazemitabaar et al., 2024; Xiao et al., 2024b). In contrast, while general-purpose LLMs are capable of pedagogical support, students often do not use them pedagogically for learning. Instead, many students take advantage of this “flexibility” to obtain direct assignment answers, leading to a false sense of competence and delayed recognition of learning gaps (Prather et al., 2024). Given that LLMs are becoming an increasingly important part of students’ educational and professional experiences, it is critical to proactively teach students how to use general-purpose LLM tools, especially AI chatbots, to facilitate their learning.
@@ 2. Related Work
2.1. Current limitations of students’ help-seeking with LLM in CS education
Help-seeking has been widely studied by educational researchers. Nelson-Le Gall ((1981)) defined help-seeking as ”the attempts to obtain assistance or intervention from another person when they cannot gratify their needs or attain a desired goal through their own efforts” and proposed a 5-stage model of the help-seeking process (Figure 1 upper-half). As a meta-cognitive skill for learning, help-seeking been strongly linked to students’ performance in CS (Marwan et al., 2020a; Bergin et al., 2005; Biró and Csernoch, 2014). Despite its importance, difficulties in seeking help effectively have been a persistent challenge in CS education (Marwan et al., 2020a) and across domains(Aleven et al., 2003). Understanding existing challenges is essential for designing instructional interventions that foster productive help-seeking behaviors in novice programmers.
Figure 1. Original model of help-seeking process (top-half) and student limitations in (LLM-Based) help-seeking (bottom-half). Blocks in Green (Step 3 and 4): Our intervention directly focuses on.
@@ 2.1. Current limitations of students’ help-seeking with LLM in CS education
Stage 2: Decision to seek help. As an additional stage not specific to helpers, LLM-related literature reported a continued, undesired pattern for both help-abuse (Kazemitabaar et al., 2023; Padiyath et al., 2024) and help-avoidance (Padiyath et al., 2024) with various reasons. For example, Kazemitabaar et al. (2023) observed CS1 students’ over-reliance (e.g. too frequent help request) on LLMs during problem-solving. Regarding help-avoidance, Padiyath et al. (2024) found a negative correlation between students’ self-reported use of LLMs in class and their concern that over-reliance on LLMs harms their programming skills. Either the over-trust or over-concerns can affect learners’ decision to seek help from LLMs.
@@
Stage 4: Employment of strategies to elicit help. Even when learners are aware of treating LLMs as tutors, they often fail to articulate their needs clearly and thus do not receive the intended responses. Nguyen et al. (2024) identified barriers such as difficulties in expressing problem and using appropriate coding terminology. Using data from a large-scale deployment of an LLM tutor in intro-level CS classrooms, Xiao et al. (2024b) found that novice CS students were incompetent in writing good prompts with a clear request and essential context (Xiao et al., 2024b), indicating students’ struggle in Stage 4 to construct the help request as perceived by the help provider (Langer and Abelson, 1972; Nelson-Le Gall, 1981).
Stage 5: Reactions to help-seeking attempt(s). When receiving responses from an LLM, novices often lack awareness or face challenges in evaluating the accuracy and reliability of the information provided. They often lack verification strategies (Amoozadeh et al., 2024), tend to over-trust the AI’s outputs (Kazemitabaar et al., 2023), and struggle to identify hallucinations, such as misleading or incorrect responses generated by the model (Xiao et al., 2024b). These can lead to the uncritical adoption of incorrect solutions, reinforcing misconceptions and hindering learning.
@@
2.2. Existing pedagogical LLM tools in CS education
@@
Over the past three years, a large number of LLM-based tutoring systems have been developed to assist novices for their programming learning (Stamper et al., 2024; Liffiton et al., 2023; Liu et al., 2024; Kazemitabaar et al., 2024; Kumar et al., 2024b; Schmucker et al., 2024; Xiao et al., 2024a; Hou et al., 2024, 2025). With extensive efforts, many of these systems iterated from a naive integration of chatbot into CS learning context, to carefully integrated learning science principles into prompt design for generating more pedagogical help. For example, there are guardrail tools that provide predefined hints or a certain type of guidance (e.g., learning-by-teaching (Schmucker et al., 2024), self-reflection (Kumar et al., 2024b), worked examples (Xiao et al., 2024a), practice problems (Hou et al., 2024)). Although incorporating learning science guidelines into predefined prompts helps prevent learners from directly copying answers, it also limits the access to adaptive support for different problems and has a weaker impact on fostering flexible help-seeking behaviors and enhancing students’ ability to seek help effectively. A limited number of studies focus on prompt engineering skills for programming, with most concentrating on crafting prompts to generate correct code (Denny et al., 2023). As a result, despite previous efforts in the pedagogical LLM-based tools, there remains a gap between supporting learners to develop more structured and pedagogically sound help-seeking behaviors when using LLMs.
@@
2.3. Link learning context to LLM-based tutoring activity: KLI Framework as the guideline
@@
With the lessons learned in the past years, the community generally agrees that when asking for LLMs for responses, directly providing solutions is an ineffective help-seeking strategy. However, what constitutes an effective prompt to learning remains underexplored (Martin-Arbos et al., 2021). In the context of CS education, while specifying that the LLM should ”behave as a programming tutor” can signal an instructional intent, it does not clarify the expected LLM response format, making the content of responses and level of pedagogy difficult to control. To address this ambiguity, it is essential not only to define the role of the LLM, but also to clearly specify the tutoring protocol, that is, the expected structure and style of the instructional interaction. Prior research, such as the Knowledge-Learning-Instruction (KLI) framework (Koedinger et al., 2012) in learning sciences suggests that teachers’ choice of instructional method should depend on the type of knowledge the student needs to acquire. The KLI framework categorizes different types of knowledge and maps them to corresponding learning events and suitable instructional methods. In this work, we use the KLI framework to guide the selection of tutoring protocols within pedagogical prompts, ensuring alignment between the knowledge type being acquired and the most effective tutoring strategies.
@@
3. Introducing Pedagogical Prompt
3.1. What is a Pedagogical Prompt?
@@
3.2. Main components of a pedagogical prompt
@@
3.2.1. Learning Context Components: A pedagogical prompt needs to be well-specified to clearly reflect the learning context
@@
3.2.2. Learner-Instructed Tutoring Protocol Component: A pedagogical prompt contains appropriate learner-instructed tutoring protocols based on learning sciences principles
@@
Figure 2. Six components, example and counter-example of a pedagogical prompt. Top: The six essential components of a pedagogical prompt, organized into Learning Context (yellow) and Tutoring Protocol (blue). Left: A complete example of a well-formed pedagogical prompt in a computing education context, including all five learning context components and a specified tutoring protocol component. Right: A counterexample with minimal context and no tutoring protocol, resulting in a less effective learning experience.
3.3. Example and counterexample implementations of a pedagogical prompt
@@
4. Formative Study: Prove the value and Understand Instructor Needs for PROOF-OF-CONCEPT Design
@@ 4.1. Demographic
Figure 3. Demographic information of 36 responded instructors: (a) countries; (b) types of degree program their institutions provide; (c) years of teaching experience in early CS courses; (d) common class size.
4.2. Confirm the need to teach pedagogical prompting
@@
4.3. Design decisions for the proof-of-concept in early undergraduate CS education
@@
4.3.1. Design-Decision-1: Instructional Method — Delivering the instruction through an interactive learning system
@@
4.3.2. Design-Decision-2: System Interaction — Including both prompt selection and prompt writing in the learning intervention
@@
4.3.3. Design-Decision-3: Length of the Training — Around 1-2 hours
@@
5. Learning Intervention: An interactive system with scenario-based instruction
@@
5.1. System Architecture and Technical Details
@@
5.2. Materials for Scenario-Based Instruction
@@
This scenario-based instruction was designed following the five key principles in Seren Smith et al. (2018). We selected four common yet unique help-seeking scenarios in early undergraduate CS learning (Hou et al., 2022; Price et al., 2017). In each scenario, the main fictional character (namely, Alice, Brian, Carol, and David; as shown in Figure 4) is facing one type of struggle during solving short programming tasks. Students are assigned the role of a peer helping a classmate (the main character). Students are required to thoroughly comprehend the scenarios and construct pedagogical prompts to guide the characters to use AI to help their learning. We gave students the autonomy to explore different prompt element combinations when creating a prompt. We also applied comic-book style narratives to help students immerse into different scenarios.
@@
Figure 4. Example scenarios designed for this study, each corresponding to a specific pedagogical prompting learning task. In the user study, Brian’s scenario is presented as the demo video, while the remaining three scenarios are shown to students during the learning session in randomized order.
@@
When learners begin the experience with our system, they are first presented with a scenario-based comic [^1] Scenario and content were designed by the authors. The comic-book style visual elements were generated by OpenAI DALL·E 3 using prompts written by the authors, then edited by authors for final use. that visually introduces the context of the prompt construction activity (Figure 5-A). The comic sets up a scenario in which learners assist fictional peers in solving programming challenges by crafting effective prompts for ChatGPT. Once students have a clear idea about the scenario and their tasks, they proceed to the learning panel, with the coding interface snapshot (including programming task description, fiction character’s current code and current output) of the given scenario on the left panel (Figure 5-B) and a step-by-step prompt builder on the right panel (Figure 5-C).
@@
Figure 5. (A) Student first views scene-setting comics for each programming problem (currently shown Alice), then the student interacts with the main learning panel, which consist of (B) left panel: a scenario snapshot including the programming task, the main character’s current code, and its output, and (C) right panel: a step-by-step prompt builder.
5.3. A Step-by-Step Pedagogical Prompt Builder

### v2

@@
Abstract
With the proliferation of large language model (LLM) applications since 2022, their use in education has sparked both excitement and concern. Recent studies consistently highlight students’ (mis)use of LLMs can hinder learning outcomes. This work aims to teach students how to effectively prompt LLMs to improve their learning. We first proposed pedagogical prompting, a theoretically-grounded new concept to elicit learning-oriented responses from LLMs. To move from concept design to a proof-of-concept learning intervention in a real educational setting, we selected early undergraduate CS education (CS1/CS2) as the example context. We began with a formative survey study with instructors teaching early-stage undergraduate-level CS courses (N=36) to inform the instructional design based on classroom needs. Based on their insights, we designed and developed a learning intervention as an interactive system with scenario-based instruction to train pedagogical prompting skills. Finally, we evaluated its instructional effectiveness through a user study with undergraduate CS novices (N=22) using pre/post-tests. Through mixed methods analyses, our results indicate significant improvements in learners’ LLM-based pedagogical help-seeking skills along with positive attitudes toward the system and increased willingness to use pedagogical prompts in their future learning. The contributions include (1) a theoretical framework of pedagogical prompting; (2) empirical insights into current instructor attitudes toward pedagogical prompting; and (3) a learning intervention design with an interactive learning tool and scenario-based instruction leading to promising results on teaching LLM-based help-seeking. Our approach is scalable for broader implementation in classrooms and has the potential to be integrated into tools like ChatGPT as an on-boarding experience to encourage learning-oriented use of generative AI.
keywords:
[…]
1 Introduction
@@
To promote more learning-oriented use of LLMs, researchers have proposed tools that embed pedagogical guardrails (Liffiton et al., 2023; Kazemitabaar et al., 2024; Xiao et al., 2024b; Hou et al., 2024). These tools aim to provide programming support without revealing direct solutions (Pirzado et al., 2024). While such designs encourage deeper engagement, they often constrain the full capabilities of LLMs and fall short of aligning with learners’ expectations or real-world needs. As a result, learners tend to underutilize these systems, favoring more versatile tools like ChatGPT, even if these are used in ways that hinder long-term learning (Kazemitabaar et al., 2024; Xiao et al., 2024a). In contrast, while general-purpose LLMs are capable of pedagogical support, students often do not use them pedagogically for learning. Instead, many students take advantage of this “flexibility” to obtain direct assignment answers, leading to a false sense of competence and delayed recognition of learning gaps (Prather et al., 2024). Given that LLMs are becoming an increasingly important part of students’ educational and professional experiences, it is critical to proactively teach students how to use general-purpose LLM tools, especially AI chatbots, to facilitate their learning.
@@ 2 Related Works
2.1 Current limitations of students’ help-seeking with LLM in CS education
Help-seeking has been widely studied by educational researchers. Nelson-Le Gall (1981) defined help-seeking as "the attempts to obtain assistance or intervention from another person when they cannot gratify their needs or attain a desired goal through their own efforts" and proposed a 5-stage model of the help-seeking process (Figure 1 upper-half). As a meta-cognitive skill for learning, help-seeking been strongly linked to students’ performance in CS (Marwan et al., 2020a; Bergin et al., 2005; Biró and Csernoch, 2014). Despite its importance, difficulties in seeking help effectively have been a persistent challenge in CS education (Marwan et al., 2020a) and across domains (Aleven et al., 2003). Understanding existing challenges is essential for designing instructional interventions that foster productive help-seeking behaviors in novice programmers.
Figure 1: Original model of help-seeking process (top-half) and student limitations in (LLM-Based) help-seeking (bottom-half). Blocks in Green (Step 3 and 4): Our intervention directly focuses on.
@@ 2.1 Current limitations of students’ help-seeking with LLM in CS education
Stage 2: Decision to seek help. As an additional stage not specific to helpers, LLM-related literature reported a continued, undesired pattern for both help-abuse (Kazemitabaar et al., 2023; Padiyath et al., 2024) and help-avoidance (Padiyath et al., 2024) with various reasons. For example, Kazemitabaar et al. (2023) observed CS1 students’ over-reliance (e.g., too frequent help request) on LLMs during problem-solving. Regarding help-avoidance, Padiyath et al. (2024) found a negative correlation between students’ self-reported use of LLMs in class and their concern that over-reliance on LLMs harms their programming skills. Either the over-trust or over-concerns can affect learners’ decision to seek help from LLMs.
@@
Stage 4: Employment of strategies to elicit help. Even when learners are aware of treating LLMs as tutors, they often fail to articulate their needs clearly and thus do not receive the intended responses. Nguyen et al. (2024) identified barriers such as difficulties in expressing problem and using appropriate coding terminology. Using data from a large-scale deployment of an LLM tutor in intro-level CS classrooms, Xiao et al. (2024a) found that novice CS students were incompetent in writing good prompts with a clear request and essential context (Xiao et al., 2024a), indicating students’ struggle in Stage 4 to construct the help request as perceived by the help provider (Langer and Abelson, 1972; Nelson-Le Gall, 1981).
Stage 5: Reactions to help-seeking attempt(s). When receiving responses from an LLM, novices often lack awareness or face challenges in evaluating the accuracy and reliability of the information provided. They often lack verification strategies (Amoozadeh et al., 2024), tend to over-trust the AI’s outputs (Kazemitabaar et al., 2023), and struggle to identify hallucinations, such as misleading or incorrect responses generated by the model (Xiao et al., 2024a). These can lead to the uncritical adoption of incorrect solutions, reinforcing misconceptions and hindering learning.
@@
2.2 Existing pedagogical LLM tools in CS education
@@
Over the past three years, a large number of LLM-based tutoring systems have been developed to assist novices for their programming learning (Stamper et al., 2024; Liffiton et al., 2023; Liu et al., 2024; Kazemitabaar et al., 2024; Kumar et al., 2024b; Schmucker et al., 2024; Xiao et al., 2024b; Hou et al., 2024, 2025). With extensive efforts, many of these systems iterated from a naive integration of chatbot into CS learning context, to carefully integrated learning science principles into prompt design for generating more pedagogical help. For example, there are guardrail tools that provide predefined hints or a certain type of guidance (e.g., learning-by-teaching (Schmucker et al., 2024), self-reflection (Kumar et al., 2024b), worked examples (Xiao et al., 2024b), practice problems (Hou et al., 2024)). Although incorporating learning science guidelines into predefined prompts helps prevent learners from directly copying answers, it also limits the access to adaptive support for different problems and has a weaker impact on fostering flexible help-seeking behaviors and enhancing students’ ability to seek help effectively. A limited number of studies focus on prompt engineering skills for programming, with most concentrating on crafting prompts to generate correct code (Denny et al., 2023). As a result, despite previous efforts in the pedagogical LLM-based tools, there remains a gap between supporting learners to develop more structured and pedagogically sound help-seeking behaviors when using LLMs.
@@
2.3 Link learning context to LLM-based tutoring activity: KLI Framework as the guideline
@@
With the lessons learned in the past years, the community generally agrees that when asking for LLMs for responses, directly providing solutions is an ineffective help-seeking strategy. However, what constitutes an effective prompt to learning remains underexplored (Martin-Arbos et al., 2021). In the context of CS education, while specifying that the LLM should "behave as a programming tutor" can signal an instructional intent, it does not clarify the expected LLM response format, making the content of responses and level of pedagogy difficult to control. To address this ambiguity, it is essential not only to define the role of the LLM, but also to clearly specify the tutoring protocol, that is, the expected structure and style of the instructional interaction. Prior research, such as the Knowledge-Learning-Instruction (KLI) framework (Koedinger et al., 2012) in learning sciences suggests that teachers’ choice of instructional method should depend on the type of knowledge the student needs to acquire. The KLI framework categorizes different types of knowledge and maps them to corresponding learning events and suitable instructional methods. In this work, we use the KLI framework to guide the selection of tutoring protocols within pedagogical prompts, ensuring alignment between the knowledge type being acquired and the most effective tutoring strategies.
@@
3 Introducing Pedagogical Prompt
3.1 What is a Pedagogical Prompt?
@@
3.2 Main components of a pedagogical prompt
@@
3.2.1 Learning Context Components: A pedagogical prompt needs to be well-specified to clearly reflect the learning context
@@
3.2.2 Learner-Instructed Tutoring Protocol Component: A pedagogical prompt contains appropriate learner-instructed tutoring protocols based on learning sciences principles
@@
Figure 2: Six components, example and counter-example of a pedagogical prompt. Top: The six essential components of a pedagogical prompt, organized into Learning Context (yellow) and Tutoring Protocol (blue). Left: A complete example of a well-formed pedagogical prompt in a computing education context, including all five learning context components and a specified tutoring protocol component. Right: A counterexample with minimal context and no tutoring protocol, resulting in a less effective learning experience.
3.3 Example and counterexample implementations of a pedagogical prompt
@@
4 Formative Study: Prove the Value and Understand Instructor Needs for PROOF-OF-CONCEPT Design
@@ 4.1 Demographic
Figure 3: Demographic information of 36 responded instructors: (a) countries; (b) types of degree program their institutions provide; (c) years of teaching experience in early CS courses; (d) common class size.
4.2 Confirm the need to teach pedagogical prompting
@@
4.3 Design decisions for the proof-of-concept in early undergraduate CS education
@@
4.3.1 Design-Decision-1: Instructional Method — Delivering the instruction through an interactive learning system
@@
4.3.2 Design-Decision-2: System Interaction — Including both prompt selection and prompt writing in the learning intervention
@@
4.3.3 Design-Decision-3: Length of the Training — Around 1-2 hours
@@
5 Learning Intervention: An Interactive System with Scenario-Based Instruction
@@
5.1 System Architecture and Technical Details
@@
5.2 Materials for Scenario-Based Instruction
@@
Figure 4: Example scenarios designed for this study, each corresponding to a specific pedagogical prompting learning task. In the user study, Brian’s scenario is presented as the demo video, while the remaining three scenarios are shown to students during the learning session in randomized order.
This scenario-based instruction was designed following the five key principles defined in the work from Seren Smith et al. (2018). We selected four common yet unique help-seeking scenarios in early undergraduate CS learning (Hou et al., 2022; Price et al., 2017). In each scenario, the main fictional character (namely, Alice, Brian, Carol, and David; as shown in Figure 4) is facing one type of struggle during solving short programming tasks. Students are assigned the role of a peer helping a classmate (the main character). Students are required to thoroughly comprehend the scenarios and construct pedagogical prompts to guide the characters to use AI to help their learning. We gave students the autonomy to explore different prompt element combinations when creating a prompt. We also applied comic-book style narratives to create immersive experience.
@@
When learners begin the experience with our system, they are first presented with a scenario-based comic [^1] The scenario and content were designed by the authors. Visual elements were generated by OpenAI DALL·E3 based on author-written prompts and subsequently edited by the authors for final use.. that visually introduces the context of the prompt construction activity (Figure 5-A). The comic sets up a scenario in which learners assist fictional peers in solving programming challenges by crafting effective prompts for ChatGPT. Once students have a clear idea about the scenario and their tasks, they proceed to the learning panel, with the coding interface snapshot (including programming task description, fiction character’s current code and current output) of the given scenario on the left panel (Figure 5-B) and a step-by-step prompt builder on the right panel (Figure 5-C).
@@
Figure 5: (A) Student first views scene-setting comics for each programming problem (currently shown Alice), then the student interacts with the main learning panel, which consist of (B) left panel: a scenario snapshot including the programming task, the main character’s current code, and its output, and (C) right panel: a step-by-step prompt builder.
5.3 A Step-by-Step Pedagogical Prompt Builder